OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form.

OpenStudy (anonymous):

y - 2 = 2x - 6 y = 2x - 4

OpenStudy (anonymous):

If you would like me to explain how I got that, i'd be happy to :P

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

So starting with y - 2 = 2(x - 3) you do the parentheses first, so 2 * x = 2x and 2 * -3 = -6

OpenStudy (anonymous):

so then you have y - 2 = 2x - 6 then to isolate the variable; y, you add + 2 to -2 to get 0 and then add 2 to the other side of the equality sign, so y = 2x - 6 + 2 so then add the negative 6 and 2 together to get -4 so y = 2x - 4

OpenStudy (anonymous):

Distribute the 2 in 2(x-3) [step 1], then add 2 to each side of the equation [step 2]: y-2=2(x-3) y-2=2x-6 [step 1] y-2 (+2)=2x-6 (+2) [step 2] y=2x-4 [simplify to get answer]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Yep :P